Liquid hexane CH 3 CH 2 4 CH 3 will react with gaseous oxygen O 2 to produce gaseous carbon dioxide CO 2 and gaseous water H 2 O . Suppose 14.6 g of hexane is mixed with 22. g of oxygen. Calculate the maximum mass of carbon dioxide that could be produced by the chemical reaction. Round your answer to 2 significant digits.
Answer #1

This reaction is combustion reaction. Balanced Combustion reaction is:

2C6H14 + 19O2 (g) 12CO2 (g) + 14H2O(g)

Mass of Hexane = 14.6 g

Molar mass of hexane = 86.18 g/mol

Moles of hexane = mass/ molar mass = 14.6 g / 86.18 g/mol = 0.169 mol

Mass of O2 = 22 g

Molar mass of O2 = 32 g/mol

Moles of O2 = 22 g / (32g/mol) = 0.6875 mol

From reaction

2 mol of Hexane reacts with 19 mol of O2

So, 0.169 mol of Hexane reacts with 19*0.169/2 = 1.61 mol of O2

Thus O2 is a limiting reagent

From reaction

19 mol of O2 produces 12 mol of CO2

0.6875 mol of O2 produces 12*0.6875/19 = 0.434 mol of CO2

Molar mass of CO2 = 44.01 g/mol

Mass of CO2 produced = moles * molar mass = 0.434 mol * 44.01 g/mol = 19.11 g = 19 g (sig. fig.)